OpenStudy (anonymous):

x power of 2 - x = 6

mathslover (mathslover):

is this your question?: \[\large{x^{2-x}=6}\] ?

mathslover (mathslover):

or this?: \[\large{x^2 - x=6}\]

OpenStudy (anonymous):

x^2 -x = 6 x^2 -x -6 = 0 (x-3)(x+2) = 0 < - solve~

OpenStudy (anonymous):

mathslover its the 2nd on you wrote.

mathslover (mathslover):

\[\large{x^2-x=6}\] \[\large{x^2-x-6=0}\] right?

OpenStudy (anonymous):

i dont understand, why would i equal it to 0?

OpenStudy (anonymous):

x^2 -x = 6 Subtract 6 from both sides x^2 -x - 6= 6 - 6 x^2 -x -6 = 0 Factor the expression (x-3)(x+2) = 0 Put the factors = 0 and solve them.

OpenStudy (vishweshshrimali5):

see \[\large \color {red} x = \color {green} 4\] subtract 4 both sides \[\large \color {red}x - 4 = \color {green} 4 - 4\] \[\large \color {red}x - 4 = 0\] similar is the case here.

OpenStudy (vishweshshrimali5):

\[\large x^2 - x = 6\] subtract 6 both sides, \[\large \color {red}{x^2 - x} - 6 = 6-6\] \[\large \color {red}{x^2 - x- 6} = \color {green} 0\]

OpenStudy (anonymous):

i understand that step, but i dont understand what comes next?

OpenStudy (vishweshshrimali5):

Now \[\large x^2 - x - 6 = 0\] \[\large x^2 - 3x + 2x - 2*3 = 0\] \[\large x(x-3) + 2(x-3) = 0\] \[\large (x+2)(x-3) = 0\] clear ?

OpenStudy (vishweshshrimali5):

@kaitlingkanatsky